OpenStudy (anonymous):

The equation is 9(v-6)=-27 solved in several steps below. For each step, choose the reason that best justifies it. 9(v-6)=-27<-given equation 9v-54=-27 9v-54+54=-27+54 9v=27 9v9=27/9 v=3 What properties are they? (multiplication of equality, subtraction of equality, distributing, simplifying, division of equality,addition of equality)

OpenStudy (texaschic101):

9(v - 6) = -27 ----given equation 9v - 54 = -27 --- distributing 9v - 54 + 54 = -27 + 54 --- addition of equality 9v = 27 -- simplifying 9v/9 = 27/9 -- division of equality v = 3

OpenStudy (anonymous):

and v=3 is simpifying right?

OpenStudy (texaschic101):

yes.....it is

OpenStudy (anonymous):

thanks again :)

OpenStudy (texaschic101):

anytime :)